# Edge Learning for 6G-enabled Internet of Things: A Comprehensive Survey of Vulnerabilities, Datasets, and Defenses

### v1

@@ Abstract
The ongoing deployment of the fifth-generation (5G) wireless networks constantly reveals limitations concerning its original concept as a key driver of Internet of Everything (IoE) applications. These 5G challenges are behind worldwide efforts to enable future networks, such as sixth-generation (6G) networks, to efficiently support sophisticated applications ranging from autonomous driving capabilities to the Metaverse. Edge learning is a new and powerful approach to training models across distributed clients while protecting the privacy of their data. This approach is expected to be embedded within future network infrastructures, including 6G, to solve challenging problems such as resource management and behavior prediction. However, edge learning in general, and distributed deep learning, in particular, have been discovered to be susceptible to tampering and manipulation. This survey article provides a holistic review of the most recent research focused on edge learning vulnerabilities and defenses for 6G-enabled IoT. We summarize the existing surveys on machine learning for 6G–IoT security and machine learning-associated threats in three different learning modes: centralized, federated, and distributed. Then, we provide an overview of enabling emerging technologies for 6G–IoT intelligence. Moreover, we provide a holistic survey of existing research on attacks against machine learning and classify threat models into eight categories, including backdoor attacks, adversarial examples, combined attacks, poisoning attacks, Sybil attacks, byzantine attacks, inference attacks, and dropping attacks. In addition, we provide a comprehensive and detailed taxonomy and a side-by-side comparison of the state-of-the-art defense methods against edge learning vulnerabilities. Finally, as new attacks and defense technologies are realized, new research and future overall prospects for 6G-enabled IoT are discussed.
@@
MEC Mobile Edge Computing
@@ I Introduction
We are experiencing accelerated development, increasing adoption, and innovative combinations of information and communication technologies, such as cloud and edge computing, the Internet of Things (IoT), massive data analytics, and Artificial Intelligence (AI). This widespread endorsement is driven by many factors, including the widespread availability of broadband communications, which are expected to move the world into an all-connected zone. One instance of such a combination is the integration of AI into the fifth-generation (5G) wireless networks. However, it is only intended to operate in specific areas under specific conditions (huge data and robust computing) [1]. This alliance is expected to be much tighter in future generations, starting with the upcoming sixth-generation (6G) wireless networks, as AI is expected to be a core component in it [2]. In addition, with Mobile Edge Computing (MEC) provides the possibility of processing large volumes of data by edge devices, and distributed learning paradigms such as Federated Learning (FL) which enable multiple parties to collaborate in building shared Machine Learning models without sharing their data. There is a lot of interest in making the edge intelligent, an emerging research area known as distributed edge learning [3].
@@
The key driver for the advancement of wireless networks has been the requirement for higher data rates, which necessitated a continuous boost in network capacity. The ongoing rise of the IoE, which is defined by Cisco as the "networked connection of people, process, data, and things" [^1] https://www.cisco.com/c/dam/en_us/about/businessinsights/docs/ioevalueindexfaq.pdf, in which billions of devices are plugged in and exchanging large amounts of data continuously, has led to a fundamental upgrade from enhanced mobile broadband (eMBB) services to ultra-reliable, low-latency communications (URLLC) [4]. While currently commercialized 5G in the ground will comfortably handle core IoE and URLLC services, as well as potentially going to support fixed access to mmWave frequencies, early 5G deployments are expected to utilize sub-6 GHz frequencies to support mobility, making it questionable whether they can provide the IoE applications of future smart cities [4]. To address these issues, 6G networks are envisioned to be the solution with AI as an indispensable component [5].
@@
Currently, not only have AI and IoT demonstrated their potential benefits in various spheres, but their synergistic effect is seen as a key factor in transforming the future, including Industry 4.0, Agriculture 4.0, and 6G communication networks. A good practical, real-world example of such cooperation is the development of research into autonomous vehicles. Furthermore, from an economic standpoint, both fields are emerging. In its 2022 AI Index report  [^2] https://aiindex.stanford.edu/wp-content/uploads/2022/03/2022-AI-Index-Report_Master.pdf, the Stanford Institute for Human-Centered Artificial Intelligence (HAI) states that investment in AI surpassed \$46 billion to reach \$93.5 billion between 2020 and 2021, with the largest growth in investment coming from the global private sector. In addition, in the same context, the worldwide cellular IoT market size is projected to expand to \$61 billion by 2026 (from \$31 billion in 2022)  [^3] https://www.juniperresearch.com/press/cellular-iot-market-value-to-exceed-61b-globally. The preceding statistics point out the success and breakthrough achieved through AI and IoT, which are virtually employed across all areas of daily lives, ranging from military, industry, healthcare, education, and entertainment, to name a few.
@@
Although both AI and IoT are considered somewhat mature from a variety of perspectives, including efficiency and operability [7], as far as their security is concerned, they are viewed as being in the early developmental stages, and further progress is required to strengthen their robustness against cyber-attacks [8, 9, 10]. To illustrate, in 2021, a widely used neural networks framework was found to have more than 300 classical security vulnerabilities, including overflows, memory corruption, bypasses, information leaks, and code executions  [^4] https://www.cvedetails.com/product/53738/Google-Tensorflow.html. A riskier pathway is the adversarial example, where core-engineered manipulations or more refined implementation of AI methods are employed to produce synthetic input data to trigger malfunctions in the targeted AI systems. As an example, a group of Skylight researchers identified a specific bias toward a specific pattern in the Cylance AI-based antivirus product, which enabled the development of a workaround by adding a selected list of strings to a malware package, altering identification scores significantly, and avoiding malware detection, with a 100% success rate for the top 10 malware (2019), and near 90% for a broader sample of 380+ malware  [^5] https://skylightcyber.com/2019/07/18/cylance-i-kill-you/. Not to mention the widespread attacks on Industrial IoT-based infrastructures in recent years, especially with sophisticated reconnaissance tools and search engines such as Shodan becoming publicly available on the Internet. We aim to shed light on these important issues within this paper (Figure 1).
@@
As we delve into cybersecurity and ML, we must recognize the interconnectedness and interdependence of two primary concepts: “security for ML” and “ML for security”. The former focuses on utilizing ML techniques to address traditional security attacks, such as denial of service (DoS), man-in-the-middle, malware, and intrusion detection. In contrast, the latter deals with the inherent vulnerabilities and potential attacks targeting ML systems, such as data poisoning. This paper explores both concepts concurrently rather than presenting them as separate entities. This approach may initially appear to create confusion; however, it is a deliberate decision driven by our belief that the two domains cannot be studied in isolation. By examining their relationship holistically, we aim to provide a comprehensive understanding of the challenges and opportunities that arise when using ML in security applications. Throughout the paper, we will demonstrate how advancements in ’security for ML’ can contribute to more robust and resilient ’ML for security’ solutions and vice versa. We believe that emphasizing the interplay between these areas is crucial for encouraging researchers and practitioners to develop innovative, integrated strategies that ensure the security and reliability of ML-based systems.
@@
Figure 2 illustrates the differentiation between Distributed Edge Learning, Federated Edge Learning, and Centralized Edge Learning. Centralized Edge Learning refers to the traditional approach to machine learning where data is collected and stored in a central location, such as an edge server or data center. In this approach, a single machine-learning model is trained using all the data available in the central location. The model is then used to make predictions on new data. On the other hand, Federated Edge Learning is a decentralized approach to machine learning where data is stored on multiple devices or servers. In this approach, the model is trained collaboratively using data from all the devices or servers without sharing the raw data. The devices or servers send model updates to a central edge server, aggregating them to form a new model. This process is repeated iteratively until the model converges to a desired level of accuracy. Distributed edge learning refers to a machine learning paradigm where a large-scale model is trained across a network of distributed devices, such as IoT devices. In this approach, each device or client is responsible for training the model using its local dataset, and then sending the updated model parameters to aggregators, which combines these updates to produce a new version of the model. Each device or client is equally privileged and forms a peer-to-peer network. Each device can communicate directly with its neighbors without a central coordinating authority. This decentralized approach is highly scalable and fault-tolerant, as it can continue to operate even if some devices fail or leave the network.
@@
The increased reliability and the enhanced utility of AI and IoT are obvious and potentially useful but they also lead to a novel and unique attack surface with cyber vulnerabilities that resemble the traditional vulnerabilities through primitive tampering and probing or a fresh category of vulnerabilities like adversarial AI. Figure 3 illustrates the effectiveness of AI and IoT as enablers of future 6G systems while focusing on the three main roles of AI: 1) operative, 2) defender, and 3) target. The main objective of this paper is to provide a comprehensive and in-depth review of threats and challenges faced by AI-based IoT systems and infrastructures. While doing so, the article focuses on the Machine Learning (ML) subset and its different learning paradigms, namely centralized, federated, and decentralized approaches. In addition, we discuss possible effective countermeasures that can be employed to protect these systems. The contributions of this study are summarized below:
@@
The structure of this article is organized as shown in Figure 1. Section II provides the review methodology. Section III sheds light on state of the art in ML in 6G-enabled IoT security and ML-associated threats. Section IV presents an overview of enabling emerging technologies for 6G–IoT intelligence. Section V discusses the use of edge learning and its application in cybersecurity. Section VI provides a detailed report on the datasets used by the scientific community. Section VII reviews the threat models attacks against machine learning and provides a classification into eight categories. Section VIII provides a taxonomy and a side-by-side comparison of the state-of-the-art defense methods. Then, we discuss new research directions and future overall prospects for 6G-enabled IoTs in Section IX. Lastly, Section X presents concluding remarks.
@@
To review and categorize the studies into two categories: ML-based security for 6G-IoT systems and ML-associated threats. The research can contribute to enhancing the security of 6G-IoT systems and identifying potential threats facing ML-associated paradigms.
Section III
@@
To provide insights into the architectural landscape, technological prospects, and security and privacy concerns associated with the use of NI in future sophisticated networks and applications.
Section IV
@@
What are the types of attacks and vulnerabilities that machine learning systems are susceptible to, particularly in 6G-IoT Networks, and how can they be classified based on the attacker’s knowledge, the type of attack employed, and the final objective?
@@
To identify and explore the challenges associated with creating reliable and trustworthy learning environments for 6G-IoT intelligence, and to propose potential solutions to these challenges.
@@
Were the potential benefits, challenges, and implications of the proposed solutions discussed in detail?
@@ II-A Objectives and research questions
In this research study, we aim to investigate the vulnerabilities, datasets, and defenses of edge learning for 6G-enabled IoT systems. Our objective is to understand the security threats posed to these systems and the methods proposed by researchers to protect them. We will analyze the vulnerabilities of machine learning and advanced deep learning techniques used to differentiate between benign and malicious traffic patterns. Our research will also explore the growing interest in defense solutions and their potential effectiveness in securing IoT systems against machine learning large-scale attacks. Table II provides the particular identified research questions to address the goals outlined above.
@@ II-B Search strategy
To identify the literature for analysis in this paper, a keyword search was conducted using terms such as "Federated Learning for 6G-enabled IoT", "Edge Learning for 6G-enabled IoT", "Distributed Learning for 6G-enabled IoT", "Federated Learning for Cyber Security", and "ML Attacks" in academic databases such as SCOPUS, Web of Science, ACM Digital Library, IEEE, Wiley, and Springer. This search produced substantial results, but some relevant primary sources may have been missed during the process. Only proposed ML attacks and Defenses methods for 6G-enabled IoT were collected, and each source was evaluated based on criteria such as reputation, relevance, originality, publication date (between $2015$ and $2023$), and influence in the field. Reputation was evaluated based on the author’s previous work and expertise in the field. The Reputation was evaluated based on the author’s previous work and expertise in the field. The Originality was evaluated based on the novelty of the ideas presented in the source. The Influence was evaluated based on the impact of the source in the field. The Sources cited more frequently were considered to have a higher influence and were given a higher score. The evaluation criteria used in this study ensured that the sources selected for analysis were of high quality, relevant to the research topic, and contained innovative and influential ideas. After evaluation, the sources were ranked based on their overall score, and the top sources were selected for analysis.
@@
Provides the title of the research paper and its citation details
@@
This describes various types of attack modes, including, centralized learning, distributed learning, federated learning
@@
White/black/Grey box
@@ II-C Data extraction
The process of data extraction is a lengthy one that involves a thorough examination of selected research papers to identify and collect essential data. To ensure that the research papers align with our research questions, a quality assessment is conducted in the initial phase, resulting in fewer papers for further analysis. The remaining papers undergo data extraction to obtain pertinent information, and Table V outlines the extracted metadata.

### v2

@@ Abstract
The deployment of the fifth-generation (5G) wireless networks in the Internet of Everything (IoE) applications and future networks (e.g., sixth-generation (6G) networks) has raised several operational challenges and limitations, for example in terms of security and privacy. Edge learning is an emerging approach to training models across distributed clients while ensuring data privacy. Such an approach when integrated in future network infrastructures (e.g., 6G) can potentially solve challenging problems such as resource management and behavior prediction. However, edge learning (including distributed deep learning) is known to be susceptible to tampering and manipulation. This survey article provides a holistic review of the extant literature focusing on edge learning-related vulnerabilities and defenses for 6G-enabled Internet of Things (IoT) systems. Existing machine learning approaches for 6G–IoT security and machine learning-associated threats are broadly categorized based on learning modes, namely: centralized, federated, and distributed. Then, we provide an overview of enabling emerging technologies for 6G–IoT intelligence. We also provide a holistic survey of existing research on attacks against machine learning and classify threat models into eight categories, namely: backdoor attacks, adversarial examples, combined attacks, poisoning attacks, Sybil attacks, byzantine attacks, inference attacks, and dropping attacks. In addition, we provide a comprehensive and detailed taxonomy and a comparative summary of the state-of-the-art defense methods against edge learning-related vulnerabilities. Finally, as new attacks and defense technologies are realized, new research and future overall prospects for 6G-enabled IoT are discussed  [^1] This paper has been accepted for publication in IEEE Communications Surveys & Tutorials: https://ieeexplore.ieee.org/document/10255264.
@@
MEC Multi-Access Edge Computing
@@ I Introduction
We are experiencing accelerated development, increasing adoption, and innovative combinations of information and communication technologies, such as cloud and edge computing, the Internet of Things (IoT), massive data analytics, and Artificial Intelligence (AI). This widespread endorsement is driven by many factors, including the widespread availability of broadband communications, which are expected to move the world into an all-connected zone. One instance of such a combination is the integration of AI into the fifth-generation (5G) wireless networks. However, it is only intended to operate in specific areas under specific conditions (massive data and robust computing) [1, 2]. This alliance is expected to be much tighter in future generations, starting with the upcoming sixth-generation (6G) wireless networks. AI is expected to be a core component in it [3, 4]. In addition, Multi-Access Edge Computing (MEC) provides the possibility of processing large volumes of data by edge devices and distributed learning paradigms such as Federated Learning (FL) which enable multiple parties to collaborate in building shared Machine Learning (ML) models without sharing their data. There is a lot of interest in making the edge intelligent, an emerging research area known as distributed edge learning [5, 6].
@@
The key driver for the advancement of wireless networks has been the requirement for higher data rates, which necessitated a continuous boost in network capacity. The ongoing rise of the Internet of Everything (IoE), which Cisco defines as the "networked connection of people, process, data, and things" [^2] https://www.cisco.com/c/dam/en_us/about/businessinsights/docs/ioevalueindexfaq.pdf, in which billions of devices are plugged in and exchanging large amounts of data continuously, has led to a fundamental upgrade from enhanced mobile broadband (eMBB) services to ultra-reliable, low-latency communications (URLLC) [7]. While currently commercialized 5G in the ground will comfortably handle core IoE and URLLC services, as well as potentially going to support fixed access to mmWave frequencies, early 5G deployments are expected to utilize sub-6 GHz frequencies to support mobility, making it questionable whether they can provide the IoE applications of future intelligent cities [7]. To address these issues, 6G networks are envisioned to be the solution with AI as an indispensable component [8].
@@
Currently, not only have AI and IoT demonstrated their potential benefits in various spheres, but their synergistic effect is seen as a critical factor in transforming the future, including Industry 4.0, Agriculture 4.0, and 6G communication networks. An excellent practical, real-world example of such cooperation is the development of research into autonomous vehicles. Furthermore, from an economic standpoint, both fields are emerging. In its 2022 AI Index report  [^3] https://aiindex.stanford.edu/wp-content/uploads/2022/03/2022-AI-Index-Report_Master.pdf, the Stanford Institute for Human-Centered Artificial Intelligence (HAI) states that investment in AI surpassed \$46 billion to reach \$93.5 billion between 2020 and 2021, with the most significant growth in investment coming from the global private sector. In addition, in the same context, the worldwide cellular IoT market size is projected to expand to \$61 billion by 2026 (from \$31 billion in 2022)  [^4] https://www.juniperresearch.com/press/cellular-iot-market-value-to-exceed-61b-globally. The preceding statistics point out the success and breakthrough achieved through AI and IoT, which are virtually employed across all areas of daily lives, ranging from military, industry, healthcare, education, and entertainment, to name a few.
@@
Although both AI and IoT are considered somewhat mature from a variety of perspectives, including efficiency and operability [10], as far as their security is concerned, they are viewed as being in the early developmental stages, and further progress is required to strengthen their robustness against cyber-attacks [11, 12, 13]. To illustrate, in 2021, a widely used neural networks framework was found to have more than 300 classical security vulnerabilities, including overflows, memory corruption, bypasses, information leaks, and code executions  [^5] https://www.cvedetails.com/product/53738/Google-Tensorflow.html. A riskier pathway is the adversarial example, where core-engineered manipulations or more refined AI methods are employed to produce synthetic input data to trigger malfunctions in the targeted AI systems. As an example, a group of Skylight researchers identified a specific bias toward a specific pattern in the Cylance AI-based antivirus product, which enabled the development of a workaround by adding a selected list of strings to a malware package, altering identification scores significantly, and avoiding malware detection, with a 100% success rate for the top 10 malware (2019), and near 90% for a broader sample of 380+ malware  [^6] https://skylightcyber.com/2019/07/18/cylance-i-kill-you/. Not to mention the widespread attacks on Industrial IoT-based infrastructures in recent years, especially with sophisticated reconnaissance tools and search engines such as Shodan becoming publicly available on the Internet. We aim to shed light on these important issues within this paper (Figure 1).
@@
As we delve into cybersecurity and ML, we must recognize the interconnectedness and interdependence of two primary concepts: “security for ML” and “ML for security”. The former focuses on utilizing ML techniques to address traditional security attacks, such as denial of service (DoS), man-in-the-middle, malware, and intrusion detection. In contrast, the latter deals with the inherent vulnerabilities and potential attacks targeting ML systems, such as data poisoning. This paper explores both concepts concurrently rather than presenting them as separate entities. This approach may initially appear to create confusion; however, it is a deliberate decision driven by our belief that the two domains cannot be studied in isolation. By examining their relationship holistically, we aim to provide a comprehensive understanding of the challenges and opportunities that arise when using ML in security applications. Throughout the paper, we will demonstrate how advancements in ’security for ML’ can contribute to more robust and resilient ’ML for security solutions and vice versa. We believe that emphasizing the interplay between these areas is crucial for encouraging researchers and practitioners to develop innovative, integrated strategies that ensure the security and reliability of ML-based systems.
@@
Figure 2 illustrates the differentiation between Distributed Edge Learning, Federated Edge Learning, and Centralized Edge Learning. Centralized Edge Learning refers to the traditional approach to machine learning where data is collected and stored in a central location, such as an edge server or data center. In this approach, a single machine-learning model is trained using all the data available in the central location. The model is then used to make predictions on new data. On the other hand, Federated Edge Learning is a decentralized approach to machine learning where data is stored on multiple devices or servers. In this approach, the model is trained collaboratively using data from all the devices or servers without sharing the raw data. The devices or servers send model updates to a central edge server, aggregating them to form a new model. This process is repeated iteratively until the model converges to a desired level of accuracy. Distributed edge learning refers to a machine learning paradigm where a large-scale model is trained across a network of distributed devices, such as IoT devices. In this approach, each device or client is responsible for training the model using its local dataset and then sending the updated model parameters to aggregators, which combines these updates to produce a new model version. Each device or client is equally privileged and forms a peer-to-peer network. Each device can communicate directly with its neighbors without a central coordinating authority. This decentralized approach is highly scalable and fault-tolerant, as it can continue to operate even if some devices fail or leave the network.
@@
The increased reliability and the enhanced utility of AI and IoT are apparent and potentially helpful. Still, they also lead to a novel and unique attack surface with cyber vulnerabilities that resemble the traditional vulnerabilities through primitive tampering and probing or a fresh category of vulnerabilities like adversarial AI. Figure 3 illustrates the effectiveness of AI and IoT as enablers of future 6G systems while focusing on the three leading roles of AI: 1) operative, 2) defender, and 3) target. The main objective of this paper is to provide a comprehensive and in-depth review of threats and challenges faced by AI-based IoT systems and infrastructures. While doing so, the article focuses on the Machine Learning (ML) subset and its different learning paradigms, namely centralized, federated, and decentralized approaches. In addition, we discuss possible effective countermeasures that can be employed to protect these systems. The contributions of this study are summarized below:
@@
The structure of this article is organized as shown in Figure 1. Section II provides the review methodology. Section IV sheds light on state of the art in ML in 6G-enabled IoT security and ML-associated threats. Section III presents an overview of enabling emerging technologies for 6G–IoT intelligence. Section V discusses the use of edge learning and its application in cybersecurity. Section VI provides a detailed report on the datasets used by the scientific community. Section VII reviews the threat models attacks against machine learning and provides a classification into eight categories. Section VIII provides a taxonomy and a side-by-side comparison of the state-of-the-art defense methods. Then, we discuss new research directions and future overall prospects for 6G-enabled IoTs in Section IX. Lastly, Section X presents concluding remarks.
@@
Guo et al. [22]
2021
No
No
No
No
No
No
No
Chen et al. [23]
2021
No
No
No
No
No
No
No
Nguyen et al. [24]
2021
Partial
Partial
No
No
No
No
No
Tang et al. [25]
2021
Partial
Partial
No
No
No
No
No
@@
Lu et al. [32]
2022
Partial
No
Partial
No
No
No
No
Vaezi et al. [33]
2022
No
No
No
No
No
No
No
Fadlullah et al. [34]
2022
Partial
Partial
No
No
No
No
Partial
Serghiou et al. [35]
2022
No
No
No
No
No
No
No
@@
To review and categorize the studies into ML-based security for 6G-IoT systems and ML-associated threats. The research can enhance the security of 6G-IoT systems and identify potential threats facing ML-associated paradigms.
Section IV
@@
To provide insights into the architectural landscape, technological prospects, and security and privacy concerns associated with using NI in future sophisticated networks and applications.
Section III
@@
What are the types of attacks and vulnerabilities are machine learning systems susceptible to, particularly in 6G-IoT Networks, and how can they be classified based on the attacker’s knowledge, the type of attack employed, and the final objective?
@@
To identify and explore the challenges associated with creating reliable and trustworthy learning environments for 6G-IoT intelligence and to propose potential solutions to these challenges.
@@
Were the proposed solutions’ potential benefits, challenges, and implications discussed in detail?
@@ II-A Objectives and research questions
In this research study, we aim to investigate the vulnerabilities, datasets, and defenses of edge learning for 6G-enabled IoT systems. We aim to understand the security threats posed to these systems and the methods researchers propose to protect them. We will analyze the vulnerabilities of machine learning and advanced deep learning techniques used to differentiate between benign and malicious traffic patterns. Our research will also explore the growing interest in defense solutions and their potential effectiveness in securing IoT systems against machine learning large-scale attacks. Table II provides the particular identified research questions to address the goals outlined above.
@@ II-B Search strategy
To identify the literature for analysis in this paper, a keyword search was conducted using terms such as "Federated Learning for 6G-enabled IoT", "Edge Learning for 6G-enabled IoT", "Distributed Learning for 6G-enabled IoT", "Federated Learning for Cyber Security", and "ML Attacks" in academic databases such as SCOPUS, Web of Science, ACM Digital Library, IEEE, Wiley, and Springer. This search produced substantial results, but some relevant primary sources may have been missed. Only proposed ML attacks and Defenses methods for 6G-enabled IoT were collected, and each source was evaluated based on criteria such as reputation, relevance, originality, publication date (between $2015$ and $2023$), and influence in the field. Reputation was evaluated based on the author’s previous work and expertise in the field. The Reputation was evaluated based on the author’s previous work and expertise in the field. The Originality was evaluated based on the novelty of the ideas presented in the source. The Influence was evaluated based on the impact of the source in the field. The Sources cited more frequently were considered to have a higher influence and were given a higher score. The evaluation criteria used in this study ensured that the sources selected for analysis were of high quality, relevant to the research topic, and contained innovative and influential ideas. After evaluation, the sources were ranked based on their overall score, and the top sources were selected for analysis.
@@
Provide the title of the research paper and its citation details
@@
This describes various types of attack modes, including centralized learning, distributed learning, federated learning
@@
White/blue/Grey box
@@ II-C Data extraction
The lengthy data extraction process thoroughly examines selected research papers to identify and collect essential data. A quality assessment is conducted in the initial phase to ensure that the research papers align with our research questions, resulting in fewer papers for further analysis. The remaining papers undergo data extraction to obtain pertinent information, and Table V outlines the extracted metadata.